OpenStudy (anonymous):

find the maclaurin series for the function: (x(5x+1))/(1-x)^3

OpenStudy (agent0smith):

http://tutorial.math.lamar.edu/Classes/CalcII/TaylorSeries_files/eq0025M.gif I know it looks complex, but you just need to find f(0), and then find a few derivatives of f

OpenStudy (anonymous):

it is odd that this function seems to have been cooked up to give all nice integer coefficients i am wondering if there is some sort of gimmick to get the answer perhaps partial fractions, but i am not sure if that is quicker

OpenStudy (anonymous):

yea i have no idea how to do it im lost

OpenStudy (anonymous):

@agent0smith do u know how to start it?

OpenStudy (agent0smith):

Take it one step at a time. First find f(0) (ie plug in 0 for x) Then find the first derivative, and f'(0) by plugging in x=0

OpenStudy (anonymous):

ok so f(0) is 0

OpenStudy (anonymous):

then i take the first deriv of the function and plug in 0 again?

OpenStudy (agent0smith):

Yes

OpenStudy (anonymous):

i was given the series for 1/(1-3)^3 does that help?

OpenStudy (anonymous):

could i take that and find the maclaurin series for the top and then just multiply it by the 1/1-x)3 answer? or no

OpenStudy (agent0smith):

Hmm, i don't think you can do that, since the top of this function is x(5x+1)... not just some constant.

OpenStudy (anonymous):

ok cause i was given that 1/(1-x)^3 = 1/2 summattion n(n-1)x^n-2 with n=2

OpenStudy (anonymous):

but idk how to use that to help me

OpenStudy (agent0smith):

I'm not sure you can... your function is more complex than just 1/(1-x)^3

OpenStudy (anonymous):

hmm im lost...

OpenStudy (anonymous):

partial fractions gives it to you if you want to use that

OpenStudy (anonymous):

too bad she is gone if you decompose in to partial fractions you get three that are easy to find the expansion of, although it is still kind of a pain, but perhaps easier than taking successive derivatives

OpenStudy (agent0smith):

That's what i considered... but it seems it's going to give us more than just the simple series... you'll have the other fractions too.

OpenStudy (anonymous):

it turns out that if you use partial fractions you get nice integer for the numerators

OpenStudy (anonymous):

of course i did it with wolfram, not by hand but you get \[\frac{5}{1-x}-\frac{11}{(1-x)^2}+\frac{6}{(1-x)^3}\] and she said she was given the expansion for the last one

OpenStudy (anonymous):

the first one is just the well known geometric series, and the second one is the derivative of that , except of course for the 11 up top then you combine them but it will still be hard to find a general formula question is rather a pain, but extra derivative will make it almost impossible quick

OpenStudy (anonymous):

@satellite73 sorry im back i had to get wifi

OpenStudy (anonymous):

oh ok, i hope you understood what i wrote it is still going to be a pain, but lots less pain than taking successive derivatives ad infinitum

OpenStudy (anonymous):

i kinda understand what u have done..what is the next step?

OpenStudy (anonymous):

compute the series for each the first one is the well known geometric series \[\frac{1}{1-x}=\sum_{n=1}^{\infty}x^k\] although here you have to multiply it by 5 and get\[\frac{5}{1-x}=5\sum_{n=1}^{\infty}x^k\]

OpenStudy (anonymous):

but what about the rest of the numerator?

OpenStudy (anonymous):

like i said, it is a pain the next one is almost \(-\frac{1}{(1-x)^2}\) which is the derivative of the first one

OpenStudy (anonymous):

the derivative of \(1+x+x^2+x^3+...\) is \(1+2x+3x^2+4x^3+...\) so it expands that way, although you still have to multiply by 11

OpenStudy (anonymous):

and you have to multiply it by \(-11\) giving \[-11\sum_{n=1}^{\infty}nx^{n-1}\]

OpenStudy (anonymous):

so far we have \[5\sum_{n=1}^{\infty}x^n-11\sum_{n=1}^{\infty}nx^{n-1}\]

OpenStudy (anonymous):

oops typo there, should start the first one at zero \[5\sum_{n=0}^{\infty}x^n-11\sum_{n=1}^{\infty}nx^{n-1}\]

OpenStudy (anonymous):

1/(1-x)^3= \[\sum_{n=2}^{\infty} n(n-1)x^n-2\]

OpenStudy (anonymous):

and lastly you said you were given the expansion for \(\frac{1}{(1-x)^3}\)

OpenStudy (anonymous):

thats what i was given

OpenStudy (anonymous):

i assume you mean \[\sum_{n=2}^{\infty}n(n-1)x^{n-2}\]

OpenStudy (anonymous):

yea i couldnt get it to go in the exponent lol

OpenStudy (anonymous):

ok so your last task it so combine them in to one series term by term i am not sure how you are going to get a nice formula out of this, but i am sure it is possible

OpenStudy (anonymous):

yea idk how to combine them into one term

OpenStudy (anonymous):

oh and you also have to remember to multiply the last one by \(-6\)

OpenStudy (anonymous):

\[5\sum_{n=0}^{\infty}x^n-11\sum_{n=1}^{\infty}nx^{n-1}-\sum_{n=2}^{\infty}n(n-1)x^{n-2}\]

OpenStudy (anonymous):

what about the 6?

OpenStudy (anonymous):

another damn typo, it is hard to write all this out \[5\sum_{n=0}^{\infty}x^n-11\sum_{n=1}^{\infty}nx^{n-1}-6\sum_{n=2}^{\infty}n(n-1)x^{n-2}\]

OpenStudy (anonymous):

its ok i am so grateful for u helping me with this

OpenStudy (anonymous):

now how about a general term? lets start with \(n=0\) and see what we get

OpenStudy (anonymous):

damn something is wrong

OpenStudy (anonymous):

oooh i see it!

OpenStudy (anonymous):

it should be \[5\sum_{n=0}^{\infty}x^n-11\sum_{n=1}^{\infty}nx^{n-1}+6\sum_{n=2}^{\infty}n(n-1)x^{n-2}\]

OpenStudy (anonymous):

now it should work better

OpenStudy (anonymous):

im lost as to what we do now? im sorry i missed the lesson on maclaurin serioes so im lost

OpenStudy (anonymous):

if \(n=0\) you get \(0\) which is what we want

OpenStudy (anonymous):

lets start with this line \[5\sum_{n=0}^{\infty}x^n-11\sum_{n=1}^{\infty}nx^{n-1}+6\sum_{n=2}^{\infty}n(n-1)x^{n-2}\]

OpenStudy (anonymous):

you want the index to start at \(n=0\) for all of them

OpenStudy (anonymous):

so we have to modify the last two correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

5∑n=0∞xn−11∑n=1∞nxn+6∑n=2∞n(n−1)xn

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

sorry idk how to use these buttons well lol

OpenStudy (anonymous):

lets change them all to \(k\) now the first one we can leave alonoe the second we want to start at 0, so we make \(k=n-1\) which makes \(n=k+1\)

OpenStudy (anonymous):

each \(n\) gets replaced by \(k+1\) and we get for the second one \[-11\sum_{k=0}^{\infty}(k+1)x^k\]

OpenStudy (anonymous):

third one we want to start at 0, so make \(k=n-2\) making \(n=k+2\) so each \(n\) gets replaced by \(k+2\)

OpenStudy (anonymous):

so then its n+2 (n+1)x^n

OpenStudy (anonymous):

oops another typo!

OpenStudy (anonymous):

\[6\sum_{k=0}^{\infty} (k+2)(k+1)x^k\]

OpenStudy (anonymous):

now they all start at the same place, and we can combine like terms as we say in the math biz

OpenStudy (anonymous):

so now we can just add them?

OpenStudy (anonymous):

\[5\sum_{n=0}^{\infty}x^n-11\sum_{k=0}^{\infty}(k+1)x^k+6\sum_{k=0}^{\infty}(k+2)(k+1)x^k\]

OpenStudy (anonymous):

yeah add them

OpenStudy (anonymous):

\[5\sum_{k=0}^{\infty}x^n-11\sum_{k=0}^{\infty}(k+1)x^k+6\sum_{k=0}^{\infty}(k+2)(k+1)x^k\]

OpenStudy (anonymous):

i guess we still have to compute \[5-11(k+1)+6(k+2)(k+1)\]

OpenStudy (anonymous):

something is bothering me here when \(k=0\) we should get \(0\)

OpenStudy (anonymous):

what happened to the x^k in the first two?

OpenStudy (anonymous):

\[5\sum_{k=0}^{\infty}x^k-11\sum_{k=0}^{\infty}(k+1)x^k+6\sum_{k=0}^{\infty}(k+2)(k+1)x^k\]

OpenStudy (anonymous):

oo so we factor out x^k right

OpenStudy (anonymous):

give me a second, let me see if i can figure out the mistake

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

damn i can't see it we should be getting 0 if \(k=0\) i think right? we know the first term of the expansion is 0

OpenStudy (anonymous):

hmm idk

OpenStudy (anonymous):

is it possible one of the coefficient s are in correct?

OpenStudy (anonymous):

idk i am not getting it right

OpenStudy (anonymous):

is works if the last coefficient is 3

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

i mean constant

OpenStudy (anonymous):

if instead of 6.. if its 3

OpenStudy (anonymous):

really? maybe we made a mistake in the expansion

OpenStudy (anonymous):

5−11(k+1)+3(k+2)(k+1)

OpenStudy (anonymous):

that gives us zero correct?

OpenStudy (anonymous):

how did u get the numbers 5 11 and 6?

OpenStudy (anonymous):

partial fraction decomposition

OpenStudy (anonymous):

from the derivative by partial fractions?

OpenStudy (anonymous):

and it ought to work, because \(5-11+6=0\)

OpenStudy (anonymous):

is that part correct?

OpenStudy (anonymous):

yes, i did it with wolfram maybe i am being stupid somewhere

OpenStudy (anonymous):

i really dont understand the steps of the maclaurin series so idk about the parts before..but i understand the changing the series to n=0 and all that looks correct to me

OpenStudy (anonymous):

lets try \(-\frac{6}{(x-1)^3}\) again we should get \(-6\) as the first term

OpenStudy (anonymous):

oh i have a idea what did you say the expansion of \(\frac{1}{(x-1)^3}\) was?

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} n(n-1)x^n-2\]

OpenStudy (anonymous):

with the -2 in the exponent

OpenStudy (anonymous):

starting at 2 right?